# Bounding the Treewidth of Outer $k$-Planar Graphs via Triangulations

### v2

@@ 3 Triangulations
We draw the vertices of $G$ on the unit circle, keeping the cyclic order defined by $\Gamma$. We set $\varepsilon>0$ to a sufficiently small value. Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices from the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Figure 2.
@@
(b) $\lambda$ and $\hat{e}$ share an endpoint; the new triangle $\triangle$ is bounded by $\lambda$, $\hat{e}$, and $c^{\prime\prime}$.
@@
Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common endpoint. In this case, we choose the other endpoint of $\hat{e}$ as $w$; see Figure 4(b). We first assume that $u$ is the common endpoint. The edge $\{u,w\}$ is an edge of $G$, so it has at most $k$ crossings. To bound the piercing number of the link $\{v,w\}$, we argue with the curve $c^{\prime\prime}$ defined in the previous case. Observe that the edge $\{u,w\}$ has exactly $j$ crossings on the part from $u$ to $x$ (including $x$). Therefore, we can argue as above and conclude that the edge $\{v,w\}$ is crossed by at most $(\ell-j)+(k-j)=k-(2j-\ell)$ edges. Note that $2j-\ell=2\lceil(\ell+1)/2\rceil-\ell$ is $1$ if $\ell$ is odd and $2$ otherwise. Similarly, if $v$ is the common endpoint, we obtain the bound $(j-1)+(k-(\ell-j+1))=k+(2j-\ell)-2$ for the piercing number of the link $\{u,w\}$.
@@
Slightly improving the analysis in the proof of Lemma 3.5 for $k=2$ yields the following.
@@
We bound the triangle piercing number more carefully. If $\lambda$ has no piercing edges or $e_{j}$ has no crossings on the right side, then the claim is clearly true. If $e_{j}$ has a crossing, the curves $c^{\prime}$ and $c^{\prime\prime}$ meet each other near $e_{j}$ with exactly $l$ crossings in total and then proceed to $w$ without crossings. Hence, the triangle piercing number is at most $4$.

### v3

@@ 3 Triangulations
We draw the vertices of $G$ on the unit circle, keeping the cyclic order defined by $\Gamma$. We set $\varepsilon>0$ to a sufficiently small value. Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices on the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Figure 2.
@@
(b) $\lambda$ and $\hat{e}$ share the endpoint $u$; the new triangle $\triangle$ is bounded by $\lambda$, $\hat{e}$, and $c^{\prime\prime}$.
@@
Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common endpoint. In this case, we choose the other endpoint of $\hat{e}$ as $w$ (see Figure 4(b)) and analyse the two subcases.
We first assume that $u$ is the common endpoint. Then the link $\{u,w\}=\hat{e}$ is an edge of $G$, so it has at most $k$ crossings. To bound the piercing number of the link $\{v,w\}$, we argue with the curve $c^{\prime\prime}$ that goes from $v$ to $w$ as defined in the previous case. Observe that the edge $\{u,w\}$ has exactly $j$ crossings on the part from $u$ to $x$ (including $x$). Therefore, we can argue as before and conclude that the piercing number of the edge $\{v,w\}$ is at most $(\ell-j)+(k-j)=k-(2j-\ell)$. This is bounded from above by $k-1$ since $2j-\ell=2\lceil(\ell+1)/2\rceil-\ell$ is $1$ if $\ell$ is odd and $2$ otherwise.
Similarly, if $v$ is the common endpoint, then the link $\{v,w\}=\hat{e}$ is an edge of $G$, so it has at most $k$ edges, and the piercing number of the link $\{u,w\}$ is at most $(j-1)+(k-(\ell-j+1))=k+(2j-\ell)-2\leq k$.
@@
With a slight modification of the strategy described in the proof of Lemma 3.5 and a careful analysis, we also obtain a better bound of the triangle piercing number for the case $k=2$.
@@
Figure 5: $\lambda$ and $\hat{e}$ share the endpoint $v$; the new triangle $\triangle$ is bounded by $\lambda$, $c^{\prime}$, and $c^{\prime\prime}$.
@@
We roughly follow the strategy of the proof of Lemma 3.5. Recall that $\lambda$ is the active link and that $\ell$ is the number of edges that pierce $\lambda$. If $\ell>0$, then $e_{j}$ with $j=\lceil(\ell+1)/2\rceil$ is the “middle” edge among those that pierce $\lambda$. If $e_{j}$ has crossings on the right side of $\lambda$, then, among these, $x$ is the crossing point of $e_{j}$ closest to $\lambda$. The edge $\hat{e}$ is the edge that crosses $e_{j}$ in $x$. As in the proof of Lemma 3.5, we distinguish the following cases.
  1. 1.
    $\lambda$ has no piercing edges.
  2. 2.
    $e_{j}$ has no crossings to the right of $\lambda$.
  3. 3.
    $\lambda$ and $\hat{e}$ are disjoint (Figure 4(a)).
  4. 4.
    $\lambda$ and $\hat{e}$ meet in $u$ (Figure 4(b)).
  5. 5.
    $\lambda$ and $\hat{e}$ meet in $v$.
In cases 14, we choose the split vertex $w$ as in the proof of Lemma 3.5, which means that the new active links are pierced at most $k=2$ times.
In case 5, we now choose $w$ to be $w_{j}$ (see Figure 5). Let $c^{\prime}$ be a curve that starts in $u$, follows $\lambda$ until just before $e_{j}$ and then follows $e_{j}$ to $w$. Let $c^{\prime\prime}$ be a curve that starts in $v$, follows $\hat{e}$ to $x$, and then follows $e_{j}$ to $w$. In order to show that each of the two curves is crossed by at most two edges, we need to do some preparations. We say that an edge lies to the left of the active link if both of its endpoints lie on the left side or one endpoint lies on the left and the other is an endpoint of the active link. We further say that an edge $e$ piercing a link $\lambda$ is anchored with respect to $\lambda$ if there exists another edge $f$ that crosses $e$ and either $f=\lambda$ or $f$ lies to the left of $\lambda$. Observe that an anchored edge can have at most one crossing to the right of $\lambda$.
We claim that, during the splitting procedure described above, every edge piercing an active link is anchored (with respect to the active link). The claim is true in the beginning as $\lambda=\{v_{1},v_{n}\}$ is not pierced. In case 1, every edge that pierces $\{u,w\}$ or $\{v,w\}$ also crosses the edge $\{u,w\}$, so it is anchored. In case 2, every edge that pierces $\{u,w\}$ or $\{v,w\}$ also pierces $\lambda=\{u,v\}$ and is anchored by induction. In case 3, observe that no edge crosses $e_{j}$ between $\lambda$ and $x$ (as $x$ is the first crossing on $e_{j}$ to the right of $\lambda$), and no edge crosses $\hat{e}$ between $x$ and $w$ (by the choice of $w$ and since $\lfloor(k-1)/2\rfloor=0$ for $k=2$). Therefore, every edge that pierces $\{u,w\}$ or $\{v,w\}$ also pierces $\lambda$ and is anchored by induction. In case 4, $\{u,w\}$ is an edge, so every edge piercing it is anchored. Every edge that pierces $\{v,w\}$ either crosses $\{u,w\}$ or pierces $\lambda$ and is anchored by induction. Similarly, in case 5, every edge that pierces $\{u,w\}$ or $\{v,w\}$ either pierces $\lambda$ or crosses $e_{j}$ (which lies to the left of both new active links).
We now return to case 5 and bound the piercing numbers of the new active links $\{u,w\}$ and $\{v,w\}$. To do this, we count the edges that cross the curves $c^{\prime}$ and $c^{\prime\prime}$, respectively. As $e_{j}$ is anchored with respect to $\lambda$, we get that $x$ is the only crossing on $e_{j}$ to the right of $\lambda$. Now, it is easy to see that $c^{\prime}$ is pierced only by $\hat{e}$ and possibly by $e_{1}$ in case $j=2$. For $c^{\prime\prime}$, we can show that there are no crossings along the curve. First, observe that any edge that crosses $\hat{e}$ between $v$ and $x$ has to pierce $\lambda$, which is not possible. Second, no edge crosses $e_{j}$ between $x$ and $w$, as $e_{j}$ is anchored and $x$ is the only crossing on $e_{j}$ to the right of $\lambda$. Combining these two observations, we see that there is no place for $c^{\prime\prime}$ to cross with an edge.
It remains to show that the triangle piercing number of the new triangle $\triangle$ formed by $u$, $v$, and $w$ is at most $4$. In case 1, $\lambda$ has no piercing edges. In cases 2, 4, and 5, $\{v,w\}$ has no piercing edges. In case 3, we distinguish two subcases depending on the relative position of $w$. When $w_{j}<w$, then $\{v,w\}$ has no piercing edges. When $w<w_{j}$, then each of $\{u,w\}$ and $\{v,w\}$ is pierced at most once.